Can you help me find the slope and y-intercept of X+Y=8 and the other problem is 2x+y=6 ?
Answer by checkley71(8403)

x+y=8 or y=x-8 then the slope is 1 and the y intercept is -8
2x+y=6 or y+2x-6 then the slope is 2 and the y intercept is -6